OpenStudy (anonymous):

How do you rewrite x^2+9 as multiplication using complex numbers and the polynomial u^2+v^2=(u+vi)(u-vi)

OpenStudy (freckles):

plug into the formula you were given

OpenStudy (freckles):

example x^2+4=(x+2i)(x-2i

OpenStudy (anonymous):

So x^2+9=(x+3i)(x+3i)?

OpenStudy (anonymous):

Wait it would be (x+3i)(x-3i)

OpenStudy (anonymous):

\[x^2+9=x^2-\left( -9 \right)=x^2-\left( 3\iota \right)^2=?\]

OpenStudy (anonymous):

So I was right?

OpenStudy (freckles):

yes (x+3i)(x-3i) is correct

OpenStudy (anonymous):

Awesome thank you!

OpenStudy (freckles):

np